OpenStudy (anonymous):

Estimate how long it will take for $1000 to double if invested at 6% compounded monthly

OpenStudy (anonymous):

\[Amount(t ( years)) = 1.06^t \times 1000\] We'll need to solve the equation: \[2000 = 1.06^t \times 1000\] \[1.06^t = { 2000 \over 1000 }\] \[1.06^t={2}\] \[^{1.06}\log ( { 2 } ) = t\] \[{ \ln(2) \over \ln(1.06) } \approx 11.9 \approx 12 \space years\]

OpenStudy (anonymous):

Here compound amount is S = 2000. principle amount is P=1000 rate of interest is r=6/100 We have to find out t The formula for S is \[\S=P(1+r)^{t}\] \[2000=1000(1+6/100)^{t}\] \[2=1.06^{t}\] log2=t log1.06 t=log2/log1.06 t=0.3010/0.0253=11.9 months = 1year (Since interest was given in months)

OpenStudy (anonymous):

ah yeah, per month. Sorry :)